OpenStudy (anonymous):

Compare the decimals 214.7 and 214.275. Which statement is true? A. 214.7 > 214.275 B. 214.7 < 214.275 C. 214.7 = 214.275

OpenStudy (anonymous):

multiply it by 10 and then compare. You will know the result. You know that multiplying by 10 shifted the decimal to right so a = 24.543*10 = 245.43, b = 24.9*10 = 249

OpenStudy (anonymous):

okay, thanks

OpenStudy (triciaal):

don't have to multiply look at the place value

OpenStudy (anonymous):

@triciaal Yes its true but I have to explain it so that he understand decimal representation.

OpenStudy (manvirk21):

A) because you have to see what number after the decimal is bigger

OpenStudy (triciaal):

just a different version if you understand the place value no need to multiply for example stating that the first place after the decimal point is the tenth place compare the digits in the corresponding locations

OpenStudy (triciaal):